Two customers,Shyam and Ekta,are visiting a particular shop in the same week (Tuesday to Saturday). Each is equally likely to visit the shop on any day as on another day. What is the probability that both will visit the shop on
$(i)$ the same day?
$(ii)$ consecutive days?
$(iii)$ different days?

Vedclass pdf generator app on play store
Vedclass iOS app on app store
(A) There are a total of $5$ days (Tuesday,Wednesday,Thursday,Friday,Saturday). Shyam can visit the shop in $5$ ways and Ekta can visit the shop in $5$ ways.
Therefore,total number of outcomes $= 5 \times 5 = 25$.
$(i)$ They can reach on the same day in $5$ ways: $(T, T), (W, W), (Th, Th), (F, F), (S, S)$.
$P(\text{same day}) = \frac{5}{25} = \frac{1}{5}$.
$(ii)$ They can reach on consecutive days in $8$ ways: $(T, W), (W, Th), (Th, F), (F, S), (W, T), (Th, W), (F, Th), (S, F)$.
$P(\text{consecutive days}) = \frac{8}{25}$.
$(iii)$ $P(\text{different days}) = 1 - P(\text{same day}) = 1 - \frac{1}{5} = \frac{4}{5}$.

Explore More

Similar Questions

$12$ defective pens are accidentally mixed with $132$ good ones. It is not possible to just look at a pen and tell whether or not it is defective. One pen is taken out at random from this lot. Determine the probability that the pen taken out is a good one.

Harpreet tosses two different coins simultaneously (say,one is of $Rs. 1$ and other of $Rs. 2$). What is the probability that she gets at least one head?

$A$ game of chance consists of spinning an arrow which comes to rest pointing at one of the numbers $1, 2, 3, 4, 5, 6, 7, 8$ (see figure),and these are equally likely outcomes. What is the probability that it will point at:
$(i)$ $8?$
$(ii)$ an odd number?
$(iii)$ a number greater than $2?$
$(iv)$ a number less than $9?$

There are $40$ students in Class $X$ of a school,of whom $25$ are girls and $15$ are boys. The class teacher has to select one student as a class representative. She writes the name of each student on a separate card,the cards being identical. Then she puts cards in a bag and stirs them thoroughly. She then draws one card from the bag. What is the probability that the name written on the card is the name of $(i)$ a girl? $(ii)$ a boy?

$A$ lot consists of $144$ ball pens of which $20$ are defective and the others are good. Nuri will buy a pen if it is good,but will not buy if it is defective. The shopkeeper draws one pen at random and gives it to her. What is the probability that
$(i)$ She will buy it?
$(ii)$ She will not buy it?

Vedclass Products

For Students

Vedclass Test Series

Mock tests in real JEE/NEET style with performance analysis. 5-day free trial.

Start Free Trial
For Teachers

Exam Paper Generator

Generate Set A/B/C/D exam papers from 7.5L+ questions in 2 minutes. 3 chapters free.

Try Free
For Institutes

Online Exam Module

Live online exams with unlimited students, 360° analytics & white-label branding.

See Demo